OpenStudy (khally92):

Now Back to CALCULUS 3 !

OpenStudy (khally92):

\[2/\pi \int\limits_{0}^{1} \int\limits_{0}^{\pi/2} \sin(u)(v+cosu)^1/2 du dv\]

OpenStudy (khally92):

@ayeshaafzal221

OpenStudy (khally92):

Team work New title Rookie. lol what's that. yay!!!.

OpenStudy (tkhunny):

Is that supposed to be \(\sqrt{v+\cos(u)}\)?

OpenStudy (khally92):

@ yes

OpenStudy (khally92):

is a problem under average value

OpenStudy (khally92):

@SolomonZelman hint?

OpenStudy (khally92):

since the region is rectangular region we can change the order of integration dvdu i think it will be much easier to integrate.

OpenStudy (solomonzelman):

when integrating with respect to v, u is constant, and vice versa

OpenStudy (tkhunny):

How did we get "u" and "v"? Have we already transformed things a bit? Try \(w = v+\cos(u)\)

OpenStudy (solomonzelman):

Latex Suggestion: When you write the power \(^{1/2}\), type `^{1/2}`

OpenStudy (khally92):

\[du=dw/-\sin(u)\] @tkhunny

OpenStudy (solomonzelman):

Also, space is `~` \(\color{#000000 }{\displaystyle 2/\pi \int\limits_{0}^{1} \int\limits_{0}^{\pi/2} \sin(u)(v+\cos u)^{1/2}~du~dv}\)

OpenStudy (khally92):

Oh nice.

OpenStudy (khally92):

kk

OpenStudy (solomonzelman):

ok, now do the sub tkhunny gave you

OpenStudy (solomonzelman):

And you should get (when integrating with respect to \(u\)); \(\color{#000000 }{\displaystyle w=v+\cos(u)}\) \(\color{#000000 }{\displaystyle -dw=\sin(u)~du}\)

OpenStudy (khally92):

@tkhunny nice one. i figured out the first integral

OpenStudy (khally92):

@SolomonZelman yes I did w=v+cos (u)

OpenStudy (khally92):

\[2/3\int\limits_{?}^{?} (v+cosu)^3/2\]

OpenStudy (tkhunny):

Rule of Thumb: ALWAYS try what's under the radical - especially in a textbook problem.

OpenStudy (khally92):

how can I tackle dv

OpenStudy (tkhunny):

Didn't learn that LaTex format, I see.

OpenStudy (khally92):

And I have pluged in my values for u \[2/3(v+\cos \pi/2)^3/2-2/3(v+1)^2/3\]

OpenStudy (solomonzelman):

\(\color{#000000 }{\displaystyle 2/\pi \int\limits_{0}^{1}\frac{-2}{3}\left[ v^{3/2}-(v+1)^{3/2}\right]~dv}\) \(\color{#000000 }{\displaystyle \frac{4}{3\pi} \int\limits_{0}^{1}\left[ (v+1)^{3/2}-v^{3/2}\right]~dv}\)

OpenStudy (solomonzelman):

this is what I got for du, and then you can integrate term by term...

OpenStudy (khally92):

\[\int\limits_{0}^{1} thats function ~~~~~~dv\]

OpenStudy (khally92):

Ok let me try

OpenStudy (khally92):

Is it different from mine /

OpenStudy (solomonzelman):

\(\color{#000000 }{\displaystyle \frac{4}{3\pi} \int\limits_{0}^{1}\left[ (v+1)^{3/2}-v^{3/2}\right]~dv}\) \(\color{#000000 }{\displaystyle \frac{4}{3\pi} \int\limits_{0}^{1} (v+1)^{3/2}dv-\frac{4}{3\pi} \int\limits_{0}^{1} v^{3/2}dv}\)

OpenStudy (solomonzelman):

what is different from what? ....

OpenStudy (khally92):

How did you get 4/3pi

OpenStudy (khally92):

instead I have \[-2/3~~~[v^3/2-(v+1)^3/2]\]

OpenStudy (solomonzelman):

\(\color{#000000 }{\displaystyle 2/\pi\int\limits_{0}^{1} \color{#0000ff}{ \int\limits_{0}^{\pi/2} \sin u(v+\cos u)^{1/2}~du}~dv}\) \(\color{#0000ff }{\displaystyle w=v+\cos u}\) \(\color{#0000ff }{\displaystyle dw=-\sin u~du\quad \Longrightarrow\quad -dw=\sin u~du}\) (I will substitute the initial variable (u), later, instead of changing the limits of integration) \(\color{#000000 }{\displaystyle 2/\pi\int\limits_{0}^{1} \color{#0000ff}{ \int\limits_{0}^{\pi/2} -(w)^{1/2}~dw}~dv}\) \(\color{#000000 }{\displaystyle 2/\pi\int\limits_{0}^{1} \color{#0000ff}{ -\left[\frac{2}{3}w^{3/2}\right]_{u=0}^{u=\pi/2}}~dv}\) \(\color{#000000 }{\displaystyle 2/\pi\int\limits_{0}^{1} \color{#0000ff}{ -\left[\frac{2}{3}(v+\cos(\pi/2))^{3/2}-\frac{2}{3}(v+\cos(0) )^{3/2}\right]}~dv}\) \(\color{#000000 }{\displaystyle 2/\pi\int\limits_{0}^{1} \color{#0000ff}{ \left[\frac{2}{3}(v+\cos(0) )^{3/2}-\frac{2}{3}(v+\cos(\pi/2))^{3/2}\right]}~dv}\) \(\color{#000000 }{\displaystyle 2/\pi\int\limits_{0}^{1} \color{#0000ff}{ \left[\frac{2}{3}(v+1 )^{3/2}-\frac{2}{3}(v+0)^{3/2}\right]}~dv}\) \(\color{#000000 }{\displaystyle \frac{4}{3\pi}\int\limits_{0}^{1} \left[(v+1 )^{3/2}-v^{3/2}\right]~dv}\) That is how I integrated that du, and notice the \(\color{red}{\displaystyle \frac{4}{3\pi}}\) in there ...

OpenStudy (solomonzelman):

Then, as I pointed out, you integrate term by term. \(\color{#000000 }{\displaystyle \frac{4}{3\pi} \int\limits_{0}^{1} (v+1)^{3/2}dv-\frac{4}{3\pi} \int\limits_{0}^{1} v^{3/2}dv}\)

OpenStudy (khally92):

I think what I have for du is right

OpenStudy (solomonzelman):

look through what I did and tell me where I made a mistake ...

OpenStudy (khally92):

LOL multiply mine by 2pi/3 and you get the same thing as yours. I multiply the constants at the end. Oh great. Nice one. I will integrate by part now.

OpenStudy (solomonzelman):

you don't mean uv-\(\int\)vdu?

OpenStudy (khally92):

the 2/pi before the integral.

OpenStudy (khally92):

No no no not by part.

OpenStudy (solomonzelman):

ok, tell me the numerical (exact value) of the integral after you integrate dv, and evaluate at the limits of integration...

OpenStudy (khally92):

ummm well I got approx 9.3084

OpenStudy (khally92):

Try again?

OpenStudy (solomonzelman):

What I get is: \(\color{#000000 }{ \displaystyle \frac{4}{3\pi} \cdot \frac{2}{5}\cdot (2)^{5/2}-\frac{4}{3\pi} \cdot \frac{2}{5}\cdot (1)^{5/2} -\frac{4}{3\pi} \cdot \frac{2}{5}\cdot (1)^{5/2}+\frac{4}{3\pi} \cdot \frac{2}{5}\cdot (0)^{5/2}}\) \(\color{#000000 }{ \displaystyle \frac{8}{15\pi} \cdot (2)^{5/2}- \frac{16}{15\pi}\cdot (1)^{5/2} }\) \(\color{#000000 }{ \displaystyle \frac{32\sqrt{2}}{15\pi} - \frac{16}{15\pi}}\) \(\color{#000000 }{ \displaystyle \frac{32\sqrt{~2~}-16}{15\pi}}\)

OpenStudy (solomonzelman):

just adding, that this is of course less than what you got...

OpenStudy (khally92):

yeah. I'll go over it again. thank you so so so so so so so so much @SolomonZelman

OpenStudy (solomonzelman):

No problem ...

OpenStudy (khally92):

Gracias:-

OpenStudy (khally92):

@SolomonZelman Please could you show me how you got \[32\sqrt{2}\]

OpenStudy (khally92):

from \[(2)^5/2\\]

OpenStudy (solomonzelman):

\(8\cdot 2^{5/2}\), actually.

OpenStudy (khally92):

yes that

OpenStudy (solomonzelman):

\(8\cdot 2^{5/2}=8\cdot2^{4/2~+~1/2}=8\cdot 2^{4/2}\cdot 2^{1/2}=8\cdot 4 \cdot \sqrt{2}=32\sqrt{2}\)

OpenStudy (solomonzelman):

that's just algebra ...

OpenStudy (solomonzelman):

I have to go for about an hour right now... hope that you have clarified everything about the problem....

OpenStudy (khally92):

Oh yea sure. thanks.